OpenStudy (anonymous):

If y is 15 times x, which equation shows the relationship between x and y?

OpenStudy (anonymous):

y = 15 + x x = 15 + y x = 15y y = 15x

OpenStudy (anonymous):

Y=15x

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Of coarse. :)